# MapReduce LoRA: Advancing the Pareto Front in Multi-Preference Optimization for Generative Models

### v2

@@ Abstract
Figure 1: MapReduce LoRA advances the Pareto fronts on Text-to-Image, Text-to-Video and language tasks. Left: On Stable Diffusion 3.5 Medium [sd35m_hf] (top) and FLUX.1-dev [flux24] (bottom), we plot the 3D Pareto front over GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25] (Column 1), and the 2D Pareto fronts where one reward weight is set to zero (Columns 2–4). Right (top): On HunyuanVideo [hunyuanvideo_arxiv'25], we plot the 2D Pareto front over Visual Quality (VQ) and Motion Quality (MQ). Right (bottom): On Llama-2 7B [llama2_arxiv23] (following Bone Soup [bonesoup_acl25]’s setup), we plot the 2D Pareto front on Helpful Assistant task over helpful and harmless rewards. Please refer to the Supplement D.1 and A for full results.
@@ 1 Introduction
Recent progress in Text-to-Image [sd3_icml24, sd35m_hf, flux24, bagel_arxiv25] and Text-to-Video [hunyuanvideo_arxiv'25, moviegen_arxiv24, wan_arxiv25] has been driven by flow-based diffusion models [flow_matching_iclr'23, rf_iclr23], achieving unprecedented visual fidelity. To better align outputs with human judgment, post-training methods, particularly reinforcement learning from human feedback (RLHF) [flow_grpo_arxiv'25, dancegrpo_arxiv25, dpo_neurips23, ppo_arxiv17], have become essential. Reward-based RLHF trains reward models from human annotations to capture specific evaluation criteria, and optimizes the generator by sampling generations, scoring them with the reward models, and updating parameters with reward-weighted objectives. However, human perception of quality is inherently multi-dimensional. In practice, multiple reward models are used to reflect criteria such as text–image alignment [GenEval_NeurIPS'23, VQAScore_ECCV24], aesthetic quality [VILA_CVPR'23], text rendering [paddleocr_arxiv'25], and overall preference [PickScore_NeurIPS'23, HPSv3_arxiv'25]. This raises a central challenge: how to optimize generative models that can improve across multiple preferences simultaneously without sacrificing any single dimension.
While post-training methods [flow_grpo_arxiv'25, dancegrpo_arxiv25] perform well under a single reward, they either degrade on unoptimized metrics at evaluation time or suffer competing gradients when jointly optimizing multiple rewards—two manifestations of the alignment tax. Multi-objective reinforcement learning (MORL) [CaPO_CVPR'25, MOPO_arxiv25, reward_soup_nips'23, selma_neurips24] attempts to optimize multiple rewards simultaneously by combining rewards through weighted mixtures or approximating a Pareto set with test-time weight control. However, weighted mixtures are dominated by easily optimized objectives, leaving harder preferences under-trained or even regressed. Calibrated Preference Optimization (CaPO) [CaPO_CVPR'25] partially addresses this imbalance but remains limited by fixed weighting and modest performance gains. Rewarded Soup [reward_soup_nips'23] offers a posteriori weight selection but still lags behind models fine-tuned individually per reward. Consequently, existing approaches struggle to deliver models that generalize well across diverse reward signals. Achieving a unified framework that improves multiple human-aligned preferences efficiently and robustly remains a fundamental challenge in multi-objective post-training.
@@
Extensive experiments demonstrate that our proposed methods substantially advance the Pareto front of multi-preference alignment across both Text-to-Image and Text-to-Video generation (Fig. 1). On Stable Diffusion 3.5 Medium [sd35m_hf] and FLUX.1-dev [flux24], our method improves GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25] by 36.1%, 4.6%, 55.7% and 32.7%, 4.3%, 67.1%, respectively. Beyond in-domain metrics, untargeted rewards—i.e., VQAScore [VQAScore_ECCV24], MPS [MPS_CVPR'24], and VILA [VILA_CVPR'23]—also improve by 1.85%, 6.49%, and 19.96%, validating the robustness and scalability of our method. On HunyuanVideo [hunyuanvideo_arxiv'25], visual and motion quality improve by 48.1% and 90.0%, achieving state-of-the-art results among post-trained diffusion systems. On language task, Helpful Assistant, with Llama-2 7B [llama2_arxiv23], helpful and harmless improve by 43.4% and 136.7%, respectively. Collectively, these results highlight the effectiveness of combining MapReduce LoRA and RaTE for efficient, controllable, and human-aligned generation. Our contributions are fourfold:
@@ 2.1 Flow-based Generative Models
Flow Matching (FM) [flow_matching_iclr'23] trains continuous normalizing flows by regressing the conditional velocity field along a path between the data distribution and a standard normal, yielding a score-free, simulation-free objective with more stable and efficient optimization than diffusion losses. Rectified Flow (RF) [rf_iclr23] specializes FM to a straight-line, approximately constant-speed path, rectifying trajectories and simplifying supervision, which improves stability and sampling efficiency. Recent Text-to-Image [sd3_icml24, sd35m_hf, flux24, bagel_arxiv25] and Text-to-Video [hunyuanvideo_arxiv'25, moviegen_arxiv24, wan_arxiv25] methods adopt FM/RF to learn the velocity field directly, improving efficiency over diffusion-style denoising and often yielding stronger gradients for conditional generation.
@@ 2.2 Reinforcement Learning from Human Feedback (RLHF)
RLHF [rlhf_neurips17, atari_rlhf_neurIPS18] was first used to train agents in simulators and Atari, and later to fine-tune language models for summarization [shf_arxiv20]. InstructGPT [instructgpt_neurips22] scaled RLHF to align broad language tasks with a three-stage pipeline: supervised fine-tuning (SFT), reward-model training, and policy optimization with Proximal Policy Optimization (PPO) [ppo_arxiv17]. Direct Preference Optimization (DPO) [dpo_neurips23] removes the reward model and online RL, directly optimizing the policy from human preference pairs. Group Relative Policy Optimization (GRPO) [grpo_arxiv'24] is a PPO-style preference method that drops the critique and uses relative scores across multiple samples per prompt, improving stability and sample efficiency. To apply RL to diffusion models, Denoising Diffusion Policy Optimization (DDPO) [ddpo_iclr24] formulates the denoising process as a Markov Decision Process (MDP). Building on this MDP formulation, Flow-GRPO [flow_grpo_arxiv'25] and DanceGRPO [dancegrpo_arxiv25] extend GRPO [grpo_arxiv'24] to flow-based generative models by using an ODE-to-SDE strategy to inject stochasticity and overcome the determinism of standard flow models.
However, single-reward RLHF optimizes one preference dimension in isolation, while generative quality spans multiple and often conflicting objectives. This creates a natural bridge to Multi-Objective RL (MORL), where the objective is to balance conflicting rewards and approximate a Pareto-optimal solution—precisely the challenge in post-training generative models. MORL methods can be categorized into two classes: a priori [CaPO_CVPR'25, MOPO_arxiv25], which scalarizes objectives into a single signal (e.g., weighted sums), and a posteriori [reward_soup_nips'23, selma_neurips24], which approximate a Pareto set and enable test-time control. CaPO [CaPO_CVPR'25] calibrates and balances multiple T2I preferences, and MOPO [MOPO_arxiv25] frames alignment as constrained, KL-regularized optimization; both are a priori and lack test-time control. Rewarded Soups [reward_soup_nips'23] extends Linear Mode Connectivity [lmc_icml'20, transfer_neurips'20] to RL, showing linear weight soups can approximate Pareto fronts and reduce reward misspecification, yet multi-reward methods often underperform specialized single-reward experts. We therefore propose two a posteriori solutions, RaTE and MapReduce LoRA, to address reward conflict and achieve comparable performances to single-reward experts’ performances.
@@ 3.1 Preliminaries
Figure 3: Overview of MapReduce LoRA and comparison with (a) individual experts, e.g., Flow-GRPO [flow_grpo_arxiv'25], and (b) Multi-Objective Reinforcement Learning, e.g., CaPO [CaPO_CVPR'25] and Rewarded soup [reward_soup_nips'23]. All methods begin from a base model $M$ and optimize with respect to reward $R$. Our proposed MapReduce LoRA iteratively trains per-reward LoRA experts and initializes iteration $k+1$ using the merged model from iteration $k$. Notably, the black dashed curve for MapReduce LoRA indicates fewer training steps compared with the black solid curves representing other methods.
Low-Rank Adaptation (LoRA) [lora_iclr'22] is a leading parameter-efficient fine-tuning method that adapts large models by learning a low-rank update to frozen pre-trained weights. For a linear layer with weight $W\in\mathbb{R}^{d\times k}$ and input $x$, the adapted layer is
@@
where $A\in\mathbb{R}^{r\times k}$, $B\in\mathbb{R}^{d\times r}$, and $r\ll\min(d,k)$. During fine-tuning, $W$ remains fixed while only $A$ and $B$ are trained, optionally scaled by a factor $\alpha/r$; at inference, $\Delta W$ can be merged into $W$. Recent work [schulman2025lora] finds that, in common post-training settings (small/medium-scale SFT, reasoning, or RL), LoRA matches full fine-tuning performance when rank and layer coverage suffice. LoRA has been used for style [multilora_tmlr24] and skill control [selma_neurips24], but multi-preference control remains underexplored; we therefore study LoRA merging for multi-preference optimization.
Textual Inversion [ti_iclr'23] learns a new embedding vector for a pseudo-token that represents a specific concept. Given only three to five reference images, it optimizes a single token embedding to encode both the high-level semantics and the fine-grained visual details of the concept, enabling personalized and reference-guided generation without modifying the base model.
Pareto Fronts (PF) [paretofront_1896] represents the collection of non-dominated solutions, model parameters for which no other candidate can improve one objective without degrading another. Formally, $\mathrm{PF}=\left\{\theta\mid\nexists\,\theta^{\prime}\in\Theta,\;\{R_{i}(\theta^{\prime})\}_{i}\succ_{\mathbb{R}^{N}}\{R_{i}(\theta)\}_{i}\right\},$ where $\succ_{\mathbb{R}^{N}}$ denotes the dominance relation in the N-dimensional reward space indexed by $i$, and $N$ denotes the total number of reward objectives.
Group Relative Policy Optimization (GRPO) [grpo_arxiv'24, flow_grpo_arxiv'25, dancegrpo_arxiv25] is a PPO-style preference optimization algorithm but based on group-normalized rewards. For each prompt $p$, GRPO samples $G$ results $\{y^{g}\}_{g=1}^{G}$ and computes group-normalized advantages $\hat{A}^{g}$ by z-scoring rewards within that group as shown in Eqn. (2). The policy is then updated with a clipped likelihood-ratio objective and regularized by a KL penalty w.r.t a frozen reference policy as shown in Eqn. (1).
@@ 3.2 MapReduce LoRA
In this section, we first demonstrate that MapReduce is a progressive souping process via averaged proximal consensus optimization. Then, we prove that MapReduce progressively converges toward the joint optimum of the conditions set by multiple reward models. Finally, we explain why MapReduce outperforms the one-shot soup method [reward_soup_nips'23]. Fig. 3 shows the overview of MapReduce LoRA and a comparison with individual experts and multi-objective reinforcement learning [CaPO_CVPR'25, reward_soup_nips'23].
Let $\{f_{i}(\theta)\}_{i=1}^{n}$ denote differentiable reward objectives, each derived from a distinct reward model used in GRPO [grpo_arxiv'24] fine-tuning. The joint optimization goal is to maximize their average:
@@ 3.3 Reward-aware Token Embedding (RaTE)
We propose RaTE to make the token embedding sensitive to user preferences and dynamically control inference-time reward influences. Inspired by Textual Inversion [ti_iclr'23], we distill each preference into a single, trainable special token embedding via supervised fine-tuning. Both teacher and student share the same frozen Transformer backbone $M$. We detail the setup below.
@@
We distill $\theta_{i}$ into $\theta_{\text{token}_{i}}$ using a Flow Matching objective [flow_matching_iclr'23]. Sample $t$ via a flow-matching timestep schedule (e.g., logit-normal or uniform (depending on the scheduler/model)), sample $\epsilon\sim\mathcal{N}(0,I)$, and set $z_{t}=(1-\sigma_{t})\,z_{0,i}^{\text{teacher}}+\sigma_{t}\,\epsilon$, where $\sigma_{t}$ is given by the scheduler. The student $M$ predicts the velocity from $z_{t}$ and the conditioned prompt $c(p,\theta_{\text{token}_{i}})$. The target velocity is $v_{\text{target}}=\epsilon-z_{0,i}^{\text{teacher}}$. The loss minimizes the Mean Squared Error (MSE) between prediction and target, updating only $\theta_{\text{token}_{i}}$:
@@ 4 Experiments
Table 1: Text-to-Image performance comparison on in-domain rewards (GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25]) within corresponding datasets and out-of-domain rewards (VQAScore [VQAScore_ECCV24], MPS [MPS_CVPR'24], and VILA [VILA_CVPR'23]) within PartiPrompts [partiprompt_TMLR22] and GenAI-Bench [GenAIBench_CVPRW24]. The performance is evaluated with fp32 precision. Red color means the performance is degraded compared to the baseline.
Method Reward GenEval [GenEval_NeurIPS'23] PickScore OCR PartiPrompts GenAIBench
@@
SD 3 M [sd3m_hf] x 0.99 0.84 0.56 0.84 0.32 0.52 0.68 - - 0.908 13.39 5.793 -
SD 3 M + CaPO [CaPO_CVPR'25] VQAScore [VQAScore_ECCV24]
+MPS [MPS_CVPR'24] 0.99 0.87 0.63 0.86 0.31 0.59 0.71 - - 0.914 13.58 5.943 -
+VILA [VILA_CVPR'23] (+0.00%) (+3.57%) (+12.50%) (+2.38%) (-3.13%) (+13.46%) (+4.41%) - - (+0.66%) (+1.42%) (+2.59%) -
SD 3.5 M [sd35m_hf] x 1.00 0.86 0.55 0.82 0.23 0.62 0.68 21.784 0.601 0.861 11.68 6.034 0.744
Individual Experts (Flow-GRPO [flow_grpo_arxiv'25]) GenEval [GenEval_NeurIPS'23] 1.00 0.96 0.93 0.91 0.93 0.84 0.93 21.852 0.654 0.879 11.81 6.101 0.782
PickScore [PickScore_NeurIPS'23] 1.00 0.94 0.76 0.88 0.36 0.59 0.76 23.359 0.716 0.880 12.66 6.755 0.774
OCR [paddleocr_arxiv'25] 1.00 0.87 0.56 0.82 0.25 0.59 0.68 21.800 0.934 0.863 11.75 6.014 0.750
Rewarded Soup [reward_soup_nips'23] [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.94 0.76 0.86 0.43 0.72 0.79 22.276 0.805 0.875 12.13 6.194 0.772
MORL-D [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.99 0.94 0.89 0.88 0.78 0.91 22.067 0.942 0.881 12.00 6.113 0.777
MORL-DR [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.82 0.61 0.81 0.21 0.59 0.67 21.834 0.641 0.860 11.71 6.014 0.746
MapReduce LoRA [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.98 0.93 0.88 0.81 0.78 0.90 22.552 0.923 0.885 12.25 6.266 0.782
@@
MapReduce LoRA + RaTE [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.98 0.93 0.90 0.95 0.79 0.92 22.777 0.936 0.877 12.44 7.238 0.777
@@
FLUX.1-dev [flux24] x 0.99 0.87 0.71 0.82 0.18 0.44 0.67 22.006 0.573 0.830 12.37 6.629 0.737
Individual Experts (Flow-GRPO [flow_grpo_arxiv'25]) GenEval [GenEval_NeurIPS'23] 1.00 1.00 0.90 0.89 0.90 0.73 0.90 22.057 0.692 0.854 12.45 6.661 0.764
PickScore [PickScore_NeurIPS'23] 1.00 0.95 0.61 0.85 0.24 0.57 0.70 23.566 0.723 0.875 12.76 5.514 0.772
OCR [paddleocr_arxiv'25] 1.00 0.84 0.76 0.83 0.20 0.46 0.68 21.997 0.971 0.841 12.32 6.591 0.749
Rewarded Soup [reward_soup_nips'23] [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 0.99 0.96 0.81 0.88 0.31 0.66 0.77 22.503 0.868 0.862 12.73 6.650 0.767
MORL-D [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.99 0.89 0.90 0.78 0.73 0.88 22.026 0.932 0.834 12.36 6.513 0.733
MORL-DR [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.95 0.90 0.94 0.73 0.79 0.88 22.112 0.952 0.840 12.41 6.534 0.745
MapReduce LoRA [GenEval_NeurIPS'23] + [PickScore_NeurIPS'23] + [paddleocr_arxiv'25] 1.00 0.96 0.88 0.88 0.81 0.79 0.89 22.951 0.957 0.875 12.95 6.572 0.775
@@
From CaPO Table 2 and 3 [CaPO_CVPR'25]; as the code is not released, we report their numbers and cannot compare CaPO on SD 3.5 M [sd35m_hf] and FLUX.1-dev [flux24].

### v3

@@ Abstract
Figure 1: MapReduce LoRA advances the Pareto fronts on Text-to-Image, Text-to-Video and language tasks. Left: On Stable Diffusion 3.5 Medium [3] (top) and FLUX.1-dev [17] (bottom), we plot the 3D Pareto front over GenEval [11], PickScore [15], and OCR [6] (Column 1), and the 2D Pareto fronts where one reward weight is set to zero (Columns 2–4). Right (top): On HunyuanVideo [16], we plot the 2D Pareto front over Visual Quality (VQ) and Motion Quality (MQ). Right (bottom): On Llama-2 7B [38] (following Bone Soup [40]’s setup), we plot the 2D Pareto front on Helpful Assistant task over helpful and harmless rewards. Please refer to the Supplement D.1 and A for full results.
@@ 1 Introduction
Recent progress in Text-to-Image [8, 3, 17, 7] and Text-to-Video [16, 31, 39] has been driven by flow-based diffusion models [22, 25], achieving unprecedented visual fidelity. To better align outputs with human judgment, post-training methods, particularly reinforcement learning from human feedback (RLHF) [23, 41, 32, 35], have become essential. Reward-based RLHF trains reward models from human annotations to capture specific evaluation criteria, and optimizes the generator by sampling generations, scoring them with the reward models, and updating parameters with reward-weighted objectives. However, human perception of quality is inherently multi-dimensional. In practice, multiple reward models are used to reflect criteria such as text–image alignment [11, 21], aesthetic quality [14], text rendering [6], and overall preference [15, 27]. This raises a central challenge: how to optimize generative models that can improve across multiple preferences simultaneously without sacrificing any single dimension.
While post-training methods [23, 41] perform well under a single reward, they either degrade on unoptimized metrics at evaluation time or suffer competing gradients when jointly optimizing multiple rewards—two manifestations of the alignment tax. Multi-objective reinforcement learning (MORL) [18, 1, 33, 20] attempts to optimize multiple rewards simultaneously by combining rewards through weighted mixtures or approximating a Pareto set with test-time weight control. However, weighted mixtures are dominated by easily optimized objectives, leaving harder preferences under-trained or even regressed. Calibrated Preference Optimization (CaPO) [18] partially addresses this imbalance but remains limited by fixed weighting and modest performance gains. Rewarded Soup [33] offers a posteriori weight selection but still lags behind models fine-tuned individually per reward. Consequently, existing approaches struggle to deliver models that generalize well across diverse reward signals. Achieving a unified framework that improves multiple human-aligned preferences efficiently and robustly remains a fundamental challenge in multi-objective post-training.
@@
Extensive experiments demonstrate that our proposed methods substantially advance the Pareto front of multi-preference alignment across both Text-to-Image and Text-to-Video generation (Fig. 1). On Stable Diffusion 3.5 Medium [3] and FLUX.1-dev [17], our method improves GenEval [11], PickScore [15], and OCR [6] by 36.1%, 4.6%, 55.7% and 32.7%, 4.3%, 67.1%, respectively. Beyond in-domain metrics, untargeted rewards—i.e., VQAScore [21], MPS [43], and VILA [14]—also improve by 1.85%, 6.49%, and 19.96%, validating the robustness and scalability of our method. On HunyuanVideo [16], visual and motion quality improve by 48.1% and 90.0%, achieving state-of-the-art results among post-trained diffusion systems. On language task, Helpful Assistant, with Llama-2 7B [38], helpful and harmless improve by 43.4% and 136.7%, respectively. Collectively, these results highlight the effectiveness of combining MapReduce LoRA and RaTE for efficient, controllable, and human-aligned generation. Our contributions are fourfold:
@@ 2.1 Flow-based Generative Models
Flow Matching (FM) [22] trains continuous normalizing flows by regressing the conditional velocity field along a path between the data distribution and a standard normal, yielding a score-free, simulation-free objective with more stable and efficient optimization than diffusion losses. Rectified Flow (RF) [25] specializes FM to a straight-line, approximately constant-speed path, rectifying trajectories and simplifying supervision, which improves stability and sampling efficiency. Recent Text-to-Image [8, 3, 17, 7] and Text-to-Video [16, 31, 39] methods adopt FM/RF to learn the velocity field directly, improving efficiency over diffusion-style denoising and often yielding stronger gradients for conditional generation.
@@ 2.2 Reinforcement Learning from Human Feedback (RLHF)
RLHF [5, 13] was first used to train agents in simulators and Atari, and later to fine-tune language models for summarization [37]. InstructGPT [29] scaled RLHF to align broad language tasks with a three-stage pipeline: supervised fine-tuning (SFT), reward-model training, and policy optimization with Proximal Policy Optimization (PPO) [35]. Direct Preference Optimization (DPO) [32] removes the reward model and online RL, directly optimizing the policy from human preference pairs. Group Relative Policy Optimization (GRPO) [36] is a PPO-style preference method that drops the critique and uses relative scores across multiple samples per prompt, improving stability and sample efficiency. To apply RL to diffusion models, Denoising Diffusion Policy Optimization (DDPO) [4] formulates the denoising process as a Markov Decision Process (MDP). Building on this MDP formulation, Flow-GRPO [23] and DanceGRPO [41] extend GRPO [36] to flow-based generative models by using an ODE-to-SDE strategy to inject stochasticity and overcome the determinism of standard flow models.
However, single-reward RLHF optimizes one preference dimension in isolation, while generative quality spans multiple and often conflicting objectives. This creates a natural bridge to Multi-Objective RL (MORL), where the objective is to balance conflicting rewards and approximate a Pareto-optimal solution—precisely the challenge in post-training generative models. MORL methods can be categorized into two classes: a priori [18, 1], which scalarizes objectives into a single signal (e.g., weighted sums), and a posteriori [33, 20], which approximate a Pareto set and enable test-time control. CaPO [18] calibrates and balances multiple T2I preferences, and MOPO [1] frames alignment as constrained, KL-regularized optimization; both are a priori and lack test-time control. Rewarded Soups [33] extends Linear Mode Connectivity [9, 28] to RL, showing linear weight soups can approximate Pareto fronts and reduce reward misspecification, yet multi-reward methods often underperform specialized single-reward experts. We therefore propose two a posteriori solutions, RaTE and MapReduce LoRA, to address reward conflict and achieve comparable performances to single-reward experts’ performances.
@@ 3.1 Preliminaries
Figure 3: Overview of MapReduce LoRA and comparison with (a) individual experts, e.g., Flow-GRPO [23], and (b) Multi-Objective Reinforcement Learning, e.g., CaPO [18] and Rewarded soup [33]. All methods begin from a base model $M$ and optimize with respect to reward $R$. Our proposed MapReduce LoRA iteratively trains per-reward LoRA experts and initializes iteration $k+1$ using the merged model from iteration $k$. Notably, the black dashed curve for MapReduce LoRA indicates fewer training steps compared with the black solid curves representing other methods.
Low-Rank Adaptation (LoRA) [12] is a leading parameter-efficient fine-tuning method that adapts large models by learning a low-rank update to frozen pre-trained weights. For a linear layer with weight $W\in\mathbb{R}^{d\times k}$ and input $x$, the adapted layer is
@@
where $A\in\mathbb{R}^{r\times k}$, $B\in\mathbb{R}^{d\times r}$, and $r\ll\min(d,k)$. During fine-tuning, $W$ remains fixed while only $A$ and $B$ are trained, optionally scaled by a factor $\alpha/r$; at inference, $\Delta W$ can be merged into $W$. Recent work [34] finds that, in common post-training settings (small/medium-scale SFT, reasoning, or RL), LoRA matches full fine-tuning performance when rank and layer coverage suffice. LoRA has been used for style [44] and skill control [20], but multi-preference control remains underexplored; we therefore study LoRA merging for multi-preference optimization.
Textual Inversion [10] learns a new embedding vector for a pseudo-token that represents a specific concept. Given only three to five reference images, it optimizes a single token embedding to encode both the high-level semantics and the fine-grained visual details of the concept, enabling personalized and reference-guided generation without modifying the base model.
Pareto Fronts (PF) [30] represents the collection of non-dominated solutions, model parameters for which no other candidate can improve one objective without degrading another. Formally, $\mathrm{PF}=\left\{\theta\mid\nexists\,\theta^{\prime}\in\Theta,\;\{R_{i}(\theta^{\prime})\}_{i}\succ_{\mathbb{R}^{N}}\{R_{i}(\theta)\}_{i}\right\},$ where $\succ_{\mathbb{R}^{N}}$ denotes the dominance relation in the N-dimensional reward space indexed by $i$, and $N$ denotes the total number of reward objectives.
Group Relative Policy Optimization (GRPO) [36, 23, 41] is a PPO-style preference optimization algorithm but based on group-normalized rewards. For each prompt $p$, GRPO samples $G$ results $\{y^{g}\}_{g=1}^{G}$ and computes group-normalized advantages $\hat{A}^{g}$ by z-scoring rewards within that group as shown in Eqn. (2). The policy is then updated with a clipped likelihood-ratio objective and regularized by a KL penalty w.r.t a frozen reference policy as shown in Eqn. (1).
@@ 3.2 MapReduce LoRA
In this section, we first demonstrate that MapReduce is a progressive souping process via averaged proximal consensus optimization. Then, we prove that MapReduce progressively converges toward the joint optimum of the conditions set by multiple reward models. Finally, we explain why MapReduce outperforms the one-shot soup method [33]. Fig. 3 shows the overview of MapReduce LoRA and a comparison with individual experts and multi-objective reinforcement learning [18, 33].
Let $\{f_{i}(\theta)\}_{i=1}^{n}$ denote differentiable reward objectives, each derived from a distinct reward model used in GRPO [36] fine-tuning. The joint optimization goal is to maximize their average:
@@ 3.3 Reward-aware Token Embedding (RaTE)
We propose RaTE to make the token embedding sensitive to user preferences and dynamically control inference-time reward influences. Inspired by Textual Inversion [10], we distill each preference into a single, trainable special token embedding via supervised fine-tuning. Both teacher and student share the same frozen Transformer backbone $M$. We detail the setup below.
@@
We distill $\theta_{i}$ into $\theta_{\text{token}_{i}}$ using a Flow Matching objective [22]. Sample $t$ via a flow-matching timestep schedule (e.g., logit-normal or uniform (depending on the scheduler/model)), sample $\epsilon\sim\mathcal{N}(0,I)$, and set $z_{t}=(1-\sigma_{t})\,z_{0,i}^{\text{teacher}}+\sigma_{t}\,\epsilon$, where $\sigma_{t}$ is given by the scheduler. The student $M$ predicts the velocity from $z_{t}$ and the conditioned prompt $c(p,\theta_{\text{token}_{i}})$. The target velocity is $v_{\text{target}}=\epsilon-z_{0,i}^{\text{teacher}}$. The loss minimizes the Mean Squared Error (MSE) between prediction and target, updating only $\theta_{\text{token}_{i}}$:
@@ 4 Experiments
Table 1: Text-to-Image performance comparison on in-domain rewards (GenEval [11], PickScore [15], and OCR [6]) within corresponding datasets and out-of-domain rewards (VQAScore [21], MPS [43], and VILA [14]) within PartiPrompts [42] and GenAI-Bench [19]. The performance is evaluated with fp32 precision. Red color means the performance is degraded compared to the baseline.
Method Reward GenEval [11] PickScore OCR PartiPrompts GenAIBench
@@
SD 3 M [2] x 0.99 0.84 0.56 0.84 0.32 0.52 0.68 - - 0.908 13.39 5.793 -
SD 3 M + CaPO [18] VQAScore [21]
+MPS [43] 0.99 0.87 0.63 0.86 0.31 0.59 0.71 - - 0.914 13.58 5.943 -
+VILA [14] (+0.00%) (+3.57%) (+12.50%) (+2.38%) (-3.13%) (+13.46%) (+4.41%) - - (+0.66%) (+1.42%) (+2.59%) -
SD 3.5 M [3] x 1.00 0.86 0.55 0.82 0.23 0.62 0.68 21.784 0.601 0.861 11.68 6.034 0.744
Individual Experts (Flow-GRPO [23]) GenEval [11] 1.00 0.96 0.93 0.91 0.93 0.84 0.93 21.852 0.654 0.879 11.81 6.101 0.782
PickScore [15] 1.00 0.94 0.76 0.88 0.36 0.59 0.76 23.359 0.716 0.880 12.66 6.755 0.774
OCR [6] 1.00 0.87 0.56 0.82 0.25 0.59 0.68 21.800 0.934 0.863 11.75 6.014 0.750
Rewarded Soup [33] [11] + [15] + [6] 1.00 0.94 0.76 0.86 0.43 0.72 0.79 22.276 0.805 0.875 12.13 6.194 0.772
MORL-D [11] + [15] + [6] 1.00 0.99 0.94 0.89 0.88 0.78 0.91 22.067 0.942 0.881 12.00 6.113 0.777
MORL-DR [11] + [15] + [6] 1.00 0.82 0.61 0.81 0.21 0.59 0.67 21.834 0.641 0.860 11.71 6.014 0.746
MapReduce LoRA [11] + [15] + [6] 1.00 0.98 0.93 0.88 0.81 0.78 0.90 22.552 0.923 0.885 12.25 6.266 0.782
@@
MapReduce LoRA + RaTE [11] + [15] + [6] 1.00 0.98 0.93 0.90 0.95 0.79 0.92 22.777 0.936 0.877 12.44 7.238 0.777
@@
FLUX.1-dev [17] x 0.99 0.87 0.71 0.82 0.18 0.44 0.67 22.006 0.573 0.830 12.37 6.629 0.737
Individual Experts (Flow-GRPO [23]) GenEval [11] 1.00 1.00 0.90 0.89 0.90 0.73 0.90 22.057 0.692 0.854 12.45 6.661 0.764
PickScore [15] 1.00 0.95 0.61 0.85 0.24 0.57 0.70 23.566 0.723 0.875 12.76 5.514 0.772
OCR [6] 1.00 0.84 0.76 0.83 0.20 0.46 0.68 21.997 0.971 0.841 12.32 6.591 0.749
Rewarded Soup [33] [11] + [15] + [6] 0.99 0.96 0.81 0.88 0.31 0.66 0.77 22.503 0.868 0.862 12.73 6.650 0.767
MORL-D [11] + [15] + [6] 1.00 0.99 0.89 0.90 0.78 0.73 0.88 22.026 0.932 0.834 12.36 6.513 0.733
MORL-DR [11] + [15] + [6] 1.00 0.95 0.90 0.94 0.73 0.79 0.88 22.112 0.952 0.840 12.41 6.534 0.745
MapReduce LoRA [11] + [15] + [6] 1.00 0.96 0.88 0.88 0.81 0.79 0.89 22.951 0.957 0.875 12.95 6.572 0.775
@@
From CaPO Table 2 and 3 [18]; as the code is not released, we report their numbers and cannot compare CaPO on SD 3.5 M [3] and FLUX.1-dev [17].